OpenStudy (anonymous):

Rewrite in standard form. Find the center and radius of the circle. (Equation below and medal goes to the best helper)

OpenStudy (anonymous):

\[2x ^{2} + 2y ^{2} - 8x + 10y + 2 = 0\]

OpenStudy (anonymous):

I have no idea how to do this :(

OpenStudy (anonymous):

center is (2,-5/2) radius is square root of 37 divided by 2

OpenStudy (anonymous):

first step is to divide by 2

OpenStudy (anonymous):

\[x^2+y^2-4x+5y+1=0\]

OpenStudy (anonymous):

can you guys draw this out ??

OpenStudy (anonymous):

next step is to complete the square twice do you know how to complete the square?

OpenStudy (anonymous):

complete the square ?? i dont believe so, i just started this lesson im a online student.

OpenStudy (anonymous):

if you do not know how to compete the square, then you cannot do it

OpenStudy (anonymous):

i can show you (maybe) it only takes a couple steps

OpenStudy (anonymous):

ok maybe your using different terminology for something I do know~ and yea that would be helpful my lessons dont help much

OpenStudy (anonymous):

ok so first off do you know what the goal is here?

OpenStudy (anonymous):

to wright the equation in standard forum

OpenStudy (anonymous):

find center and radius

OpenStudy (anonymous):

right, which is \[(x-h)^2+(y-k)^2=r^2\] notice that \((x-h)^2\) and \(y-k)^2\) are "perfect squares" i.e. the square of something that is why you have to "complete the square" twice once for the \(x\) terms and once for the \(y\) terms

OpenStudy (anonymous):

lets group them together first \[x^2-4x+y^2+5y=-1\]

OpenStudy (anonymous):

oh thats what u meant never heard that term ~

OpenStudy (anonymous):

so we have to fill in these blanks for \(h\) and \(k\) \[(x-h)^2+(y-k)^2=-1+something\]

OpenStudy (anonymous):

what is half of 4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

and what is \(2^2\)?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so our start is \[(x-2)^2+(y-k)^2=-1+4=3\] now we repeat for the y term

OpenStudy (anonymous):

what is half of 5?

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

yes, but don't use a decimal

OpenStudy (anonymous):

so what do i use ? 2 ? 3?

OpenStudy (anonymous):

no, a fraction

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 }\] ?

OpenStudy (anonymous):

you got it upside down!!

OpenStudy (anonymous):

oh soo \[\frac{ 5 }{ 2}\]

OpenStudy (anonymous):

yeah and what is \((\frac{5}{2})^2\)?

OpenStudy (anonymous):

mmm gosh multiplying fractions :( ummm is it ... \[\frac{ 25 }{ 2 }\]

OpenStudy (anonymous):

no, you forgot to square the two

OpenStudy (anonymous):

ohh so \[\frac{ 25 }{ 4 }\]

OpenStudy (anonymous):

I always thought you left the bottom number alone

OpenStudy (anonymous):

right so now we have \[(x-2)^2+(y+\frac{5}{2})^2=3+\frac{25}{4}\]

OpenStudy (anonymous):

that is called "completing the square" you have two competed squares on the left almost done, last step is only to add on the right, then it is in standard form

OpenStudy (anonymous):

you leave the bottom number alone if you are ADDING fractions with like denominators for example \[\frac{5}{2}=\frac{8}{2}=\frac{13}{2}\]

OpenStudy (anonymous):

but if you are multiplying you just multiply \[\left(\frac{5}{2}\right)^2=\frac{5}{2}\times \frac{5}{2}=\frac{25}{4}\]

OpenStudy (anonymous):

ohh so this will be \[\left( x - 2 \right)^{2} +\left( y +\frac{ 5 }{ 2 } \right)^{2} = \frac{ 25 }{ 4} 3\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what the heck kind of number is \(\frac{ 25 }{ 4} 3\)?? never seen a number that looked like that before

OpenStudy (anonymous):

you gotta add !

OpenStudy (anonymous):

well I thought i was adding I guess not ! ~ -.-

OpenStudy (anonymous):

the way you wrote it, it looks like some sort of multiplication not sure in any case you get a fraction when you add

OpenStudy (anonymous):

I added the 3 to the 25/4 at the end ~

OpenStudy (anonymous):

\[3+\frac{25}{4}=\frac{3\times 4+25}{4}=\frac{37}{4}\]

OpenStudy (anonymous):

oh I just added it as an whole number .. nvm alright soo i add that to the end and its the answer ?

OpenStudy (anonymous):

ok well what about center and radius ?

OpenStudy (anonymous):

\[(x-2)^2+(y+\frac{5}{2})^2=\frac{37}{4}\] is standard from you read the center and radius from that form

OpenStudy (anonymous):

the center is \[(2,-\frac{5}{2})\] and the radius is \[\frac{\sqrt{37}}{2}\]

OpenStudy (anonymous):

ohhhh ~~~ how so they where already in the equation ?